OpenStudy (anonymous):

Write each expression as a single logarithm 2 log(x) + 3 log(x+1)

mathslover (mathslover):

a log(x) can be written as \(\log(x^a)\) So, 2 \(\log(x)\) = ?

OpenStudy (anonymous):

log(x^2)?

mathslover (mathslover):

Right. Similarly do with 3 log(x+1)

OpenStudy (anonymous):

3 log(x+1) = log(x+1)^3

mathslover (mathslover):

Very good! So, you get : \(2 \log x + 3 \log(x+1) = \log (x^2) + \log (x+1)^3 \) Now, use : log a +log b = log (ab)

OpenStudy (anonymous):

log(x^2) * log(x+1)^3

mathslover (mathslover):

No, it you have to put a and b in single log. That is : \(\log ( (x^2) \times (x+1)^3 ) \)

OpenStudy (anonymous):

can that be simplified?

mathslover (mathslover):

Well, you can expand (x+1)^3 and then multiply x^2 and it .

mathslover (mathslover):

But, I think, if you put like the above expression only, it will be fine too!

OpenStudy (anonymous):

Thanks, I have one more to post so I will close this question

mathslover (mathslover):

Sure. Tag me there, I will try to help.

OpenStudy (anonymous):

Ok I will do that :)